# Rheo-SINDy: Finding a constitutive model from rheological data for complex fluids using sparse identification for nonlinear dynamics

### v2

@@ III Case Studies
The dumbbell-based models have been widely utilized in numerous previous studies for the computation of viscoelastic fluids and are regarded as the standard mesoscopic model for viscoelastic fluids Bird, Armstrong, and Hassager (1987). As illustrated in Fig. 2, a dumbbell consists of two beads (indexed as $1$ or $2$) and a spring that connects them.
@@
where $\langle R_{\rm eq}^{2}\rangle^{1/2}=(n_{\rm K})^{1/2}b_{\rm K}$ is the equilibrium length of the springs, and $R_{\rm max}=n_{\rm K}b_{\rm K}$ is the maximum length of the springs. As shown later in Sec. III.3, a constitutive equation cannot be analytically obtained for the FENE dumbbell model. To address the FENE spring more analytically, the following approximate expression of the FENE spring has been proposed Bird, Armstrong, and Hassager (1987):
@@
III.1 Hookean Dumbbell Model
The most basic dumbbell model is the Hookean dumbbell model, where Hookean springs are employed (cf. Eq. (16)). From Eqs. (12), (15), and (16), the Hookean dumbbell model reduces to the constitutive equation for the UCM model (cf. Eq. (S5) in the supplementary material) in the limit of $N_{\rm p}\rightarrow\infty$ with $N_{\rm p}$ being the number of dumbbells. Specifically, the constitutive equations of the UCM model for each stress component under shear flow are expressed as
${\dot{\tau}}_{xx}=-\tau_{xx}+2\tau_{xy}\kappa_{xy},$ (19)
${\dot{\tau}}_{yy/zz}={\tau}_{yy/zz}=0,$ (20)
Here, “$yy/zz$” in Eq. (20) means “$yy$ or $zz$”.
For the Hookean dumbbell model, we generate training data by Brownian dynamics (BD) simulations with the finite numbers of dumbbells ($N_{\rm p}\in\{10^{3},10^{4},10^{5}\}$). We apply the oscillatory shear flow, $\gamma(t)=\gamma_{0}\sin(\omega t)$ (i.e., $\kappa_{xy}(t)=\dot{\gamma}(t)=\gamma_{0}\omega\cos(\omega t)$), with $\gamma_{0}=2$ and $\omega=0.5$, over a period from $t=0$ to $t=100$. The simulations are run with $\Delta t=1\times 10^{-3}$ for $0\leq t\leq 100$ and data are collected at the interval of $\Delta t_{\rm train}=1\times 10^{-2}$. Each simulation is conducted with five different random seeds, and their average data is used for training. Due to the characteristics of the BD simulation, the training data inherently include noise originating from the finite $N_{\rm p}$. We here test whether Rheo-SINDy can find from the noisy data the constitutive equations for the UCM model shown in Eq. (19)–(21).
III.2 FENE-P Dumbbell Model
@@
$\frac{{\rm d}{\bm{C}}}{{\rm d}t}-{\bm{C}}\cdot{\bm{\kappa}}^{+}-{\bm{\kappa}}\cdot{\bm{C}}=-f_{\rm FENE}(t){\bm{C}}+\frac{n_{\rm K}}{3}{\bm{I}},$ (22)
where $\bm{\kappa}^{+}$ is the transposed $\bm{\kappa}$. The stress tensor is thus obtained by
@@
$\frac{{\rm d}{\bm{\tau}}}{{\rm d}t}-{\bm{\tau}}\cdot{\bm{\kappa}}^{+}-{\bm{\kappa}}\cdot{\bm{\tau}}=-f_{\rm FENE}(t){\bm{\tau}}+2{\bm{D}}+\frac{{\rm D}\ln Z}{{\rm D}t}({\bm{\tau}}+{\bm{I}}),$ (27)
@@
To generate noise-free training data, we solve Eqs. (23)–(26) with $n_{\rm K}=10$ and $\Delta t=1\times 10^{-4}$ for $0\leq t\leq 100$. We apply the oscillatory shear flow with $\gamma_{0}=2$ and various $\omega$ values ($\omega\in\{0.1,0.2,\ldots,1\}$). From the computed stress data, we collect data at the interval of $\Delta t_{\rm train}=1\times 10^{-2}$.
III.3 FENE Dumbbell Model
@@ IV Results and Discussions
In this section, we present the results of the case studies for the dumbbell models. From the case studies on phenomenological constitutive equations shown in Sec. S3 in the supplementary material, we have made the following two findings: (i) taking shear rheological data by an oscillatory shear test is more appropriate than by a simple (constant) shear test, and (ii) among the five optimization methods shown in Sec. II, the STRidge or a-Lasso is superior to the other three methods. Thus, in what follows, we generate the training data using the oscillatory shear test and employ the STRidge and a-Lasso as optimization methods. To discuss the effect of the hyperparameter $\alpha$, we also examine the results with $\alpha$, not selected in the proposed method.
IV.1 Hookean Dumbbell Model
Figure 3: (a) The total number of terms and (b) the constitutive equations obtained by Rheo-SINDy with the STRidge (black) and a-Lasso (red) for the Hookean dumbbell model (Eqs. (19)–(21)). In (a), the horizontal line indicates the correct number of terms, and circle, triangle, and square symbols represent the results for $N_{\rm p}=10^{3}$, $10^{4}$, and $10^{5}$, respectively.
We first explain the results for the Hookean dumbbell model. We here used the polynomial library that includes up to second order terms of $\tau_{xx}$, $\tau_{yy}$, $\tau_{xy}$, and $\kappa_{xy}$. Thus, the total number of candidate terms was $N_{\bm{\Theta}}=15$ for each component.
Figure 3 shows the Rheo-SINDy results for the Hookean dumbbell model with the different numbers of dumbbells. We note that the standard deviation of $\bm{\tau}$ in the training data decreases proportionally with $N_{\rm p}^{-1/2}$. From Fig. 3(a), as the value of $N_{\rm p}$ increases, sparser solutions are obtained, especially for Rheo-SINDy with the STRidge. Unlike the case of the UCM model (cf. Fig. S1 in the supplementary material), which can be considered as the “noise-free” case of the Hookean dumbbell model, the STRidge provides the correct number of terms only within a narrow range of $\alpha$ values. Nevertheless, if we choose the appropriate $\alpha$ value, the (nearly) correct constitutive equations can be found by the STRidge, as shown in the upper part of Fig. 3(b). We note that the terms containing $\tau_{yy}$ appear in the time evolution equation for $\tau_{xy}$ obtained by the STRidge. Although these terms do not affect the predictions because ${\tau}_{yy}=0$, these terms do not appear in the correct equation. We speculate that the appearance of these terms is due to the correlation effects of the noise in $R_{x}$ and $R_{y}$ on the stress (cf. Eq. (15)). When comparing the STRidge and a-Lasso, it is evident that the a-Lasso provides stable and sparse solutions across a broader range of $\alpha$ values, regardless of the $N_{\rm p}$ value. Furthermore, we confirm that the correct equations can be obtained using the a-Lasso, as shown in the lower part of Fig. 3(b). This partially suggests the effectiveness of the a-Lasso in discovering essential terms from noisy data.
Figure 4: (a) The total number of terms and (b) the error rate of the constitutive equations obtained by Rheo-SINDy with the STRidge (black squares) and a-Lasso (red reverse triangles) for the FENE-P dumbbell model in conformation expression (Eqs. (24)–(26) with $n_{\rm K}=10$). The horizontal line in (a) indicates the correct number of terms. The blue circles represent the $\alpha$ values selected for test simulations.
Figure 5: The constitutive equations obtained by Rheo-SINDy with the STRidge and a-Lasso for the FENE-P dumbbell model in conformation expression (Eqs. (24)–(26) with $n_{\rm K}=10$).
@@ IV.1 Hookean Dumbbell Model
Figure 6: Test simulation results under the oscillatory shear flow with $\gamma_{0}=4$ and $\omega=1$ using the equations obtained by Rheo-SINDy with (a) the STRidge and (b) a-Lasso for the FENE-P dumbbell model in conformation expression. The obtained data from the equations were converted to the stress using the dimensionless form of Eq. (23). The black, blue, and red lines show $\tau_{xx}$, $\tau_{yy}$, and $\tau_{xy}$. The bold, thin dotted, and thin solid lines indicate the exact solutions (Eqs. (23)–(26) with $n_{\rm K}=10$), predictions with smaller $\alpha$ values ($\alpha=1\times 10^{-3}$ for the STRidge and $\alpha=1\times 10^{-7}$ for the a-Lasso), and predictions with larger $\alpha$ values ($\alpha=1\times 10^{-1}$ for the STRidge and $\alpha=1\times 10^{-4}$ for the a-Lasso).
IV.2 FENE-P Dumbbell Model
@@
We next examine whether Rheo-SINDy can find more complex differential equations (i.e., the FENE-P dumbbell model) than the UCM model and the Giesekus model. We utilized Rheo-SINDy with $\bm{T}$ replaced by $\bm{C}$ to discover the differential equations for the conformation tensor $\bm{C}$ of the FENE-P dumbbell model explained in Sec. III.2. In this case, we prepared the following custom library:
@@
Figure 4 displays the results: (a) the total number of predicted terms and (b) the error rate as a function of the hyperparameter $\alpha$. The error rate is defined as the sum of the mean squared errors (MSEs) of $\dot{\bm{t}}_{\mu\nu}-{\bm{\Theta}}\hat{\bm{\xi}}_{\mu\nu}$ scaled so that the maximum value of each method is 1. Similar to the results for the phenomenological constitutive equations shown in Sec. S3 in the supplementary material, the a-Lasso provides sparser solutions than the STRidge, and the STRidge gives lower error rates than the a-Lasso. Figure 5 presents the differential equations obtained by the STRidge and a-Lasso for two $\alpha$ values that yield the nearly correct number of terms with a small error rate. From the lower part of Fig. 5, while the a-Lasso can provide sparser solutions, they are not correct (cf. Eqs. (24)–(26)). Specifically, in all cases for $C_{xx}$, $C_{yy}$, and $C_{zz}$, the a-Lasso failed to identify the constant term in Eqs. (24) and (25), which is a possible source of larger errors compared to the STRidge. In the case of the STRidge, we confirmed that by choosing the appropriate $\alpha$ ($\alpha=1\times 10^{-1}$), nearly correct differential equations can be obtained, as shown in the upper part of Fig. 5. Since the $yy$-component and $zz$-component of the stress are equivalent, the exact equations can be recovered by setting $C_{yy}=C_{zz}$. Thus, we conclude that the correct differential equations for the FENE-P dumbbell model can be obtained by Rheo-SINDy if we properly prepare the library functions and tune the hyperparameter.
@@
To validate the obtained equations in Fig. 5, we generated data using the equations under conditions different from those used to generate the training data (namely, the oscillatory shear flow with $\gamma_{0}=4$ and $\omega=1$), and converted these data to the stress data using the dimensionless form of Eq. (23). Figure 6 compares the generated data with the correct data generated from Eqs. (23)–(26). As shown in this figure, the equations obtained by the STRidge can reproduce the exact solutions even when the equations are not exactly correct ($\alpha=1\times 10^{-3}$). In contrast, the test simulations with the differential equations obtained by the a-Lasso show the deviations from the test data, especially for $\tau_{xx}$. These results emphasize the need to choose an appropriate optimization method to obtain reasonable solutions.
@@
Figure 7: (a) The total number of terms and (b) the error rate of the constitutive equation obtained by Rheo-SINDy with the STRidge (black squares) and the a-Lasso (red reverse triangles) for the FENE-P dumbbell model in stress expression (Eqs. (29)–(31) with $n_{\rm K}=10$). The horizontal short-dashed line in (a) indicates that the number of terms is zero. The blue circles represent the $\alpha$ values selected for test simulations.
Figure 8: The constitutive equations obtained by Rheo-SINDy with the STRidge and a-Lasso for the FENE-P dumbbell model in stress expression (Eqs. (29)–(31) with $n_{\rm K}=10$).
@@
Figure 9: Test simulation results under the oscillatory shear flow with $\gamma_{0}=4$ and $\omega=1$ using the constitutive equations obtained by Rheo-SINDy with (a) the STRidge and (b) a-Lasso for the FENE-P dumbbell model in stress expression. The black, blue, and red lines represent the $xx$-, $yy$-, and $xy$-components of the stress tensor, respectively. The bold lines show the exact solutions. The thin solid and short-dashed lines indicate the results with smaller $\alpha$ values ($\alpha=1\times 10^{-3}$ for the STRidge and $\alpha=1\times 10^{-8}$ for the a-Lasso) and with larger $\alpha$ values ($\alpha=1$ for the STRidge and $\alpha=1\times 10^{-4}$ for the a-Lasso).

### v3

@@ III.3 Dumbbell Models
The dumbbell-based models have been widely utilized in numerous previous studies for the computation of viscoelastic fluids and are regarded as the standard mesoscopic model for viscoelastic fluids Bird, Armstrong, and Hassager (1987). As a stepping stone to investigating more complex models, we chose this model, which can be considered as a fundamental model in rheology. As illustrated in Fig. 2, a dumbbell consists of two beads (indexed as $1$ or $2$) and a spring that connects them.
@@
where $\langle R_{\rm eq}^{2}\rangle^{1/2}=(n_{\rm K})^{1/2}b_{\rm K}$ is the equilibrium length of the springs, and $R_{\rm max}=n_{\rm K}b_{\rm K}$ is the maximum length of the springs. As shown later in Sec. III.3.3, a constitutive equation cannot be analytically obtained for the FENE dumbbell model. To address the FENE spring more analytically, the following approximate expression of the FENE spring has been proposed Bird, Armstrong, and Hassager (1987):
@@
III.3.1 Hookean Dumbbell Model
The most basic dumbbell model is the Hookean dumbbell model, where Hookean springs are employed (cf. Eq. (32)). From Eqs. (28), (31), and (32), the Hookean dumbbell model reduces to the constitutive equation for the UCM model (cf. Eq. (18)) in the limit of $N_{\rm p}\rightarrow\infty$ with $N_{\rm p}$ being the number of dumbbells.
For the Hookean dumbbell model, we generate training data by Brownian dynamics (BD) simulations with the finite numbers of dumbbells ($N_{\rm p}\in\{10^{3},10^{4},10^{5}\}$). In the BD simulations of this study, to integrate Eq. (28), we use the explicit Euler method with a small time step. We apply the oscillatory shear flow, $\gamma(t)=\gamma_{0}\sin(\omega t)$, with $\gamma_{0}=2$ and $\omega=0.5$, over a period from $t=0$ to $t=100$. The simulations are run with $\Delta t=1\times 10^{-3}$ for $0\leq t\leq 100$ and data are collected at the interval of $\Delta t_{\rm train}=1\times 10^{-2}$. Each simulation is conducted with five different random seeds, and their average data is used for training. Due to the characteristics of the BD simulation, the training data inherently include noise originating from the finite $N_{\rm p}$. We here test whether Rheo-SINDy can find from the noisy data the constitutive equations for the UCM model shown in Eqs. (19)–(21).
Table 3: Settings of Rheo-SINDy used in case studies.
model exact equations optimization [^6] The value of $\alpha$ is picked from sparser solutions whose error is the same order of magnitude as the minimum error. library $N_{\bm{\Theta}}$
UCM Eqs. (19)–(21) all five methods [^7] All five methods are the STLSQ, STRidge, Lasso, E-Net, and a-Lasso explained in Sec. II. polynomial (up to 3rd order) 35
Giesekus Eqs. (23)–(26) STLSQ, STRidge, a-Lasso polynomial (up to 2nd order) 15
Hookean dumbbell Eqs. (19)–(21) STRidge, a-Lasso polynomial (up to 2nd order) 15
FENE-P dumbbell Eqs. (37)–(39) STRidge, a-Lasso Eq. (45) 26
FENE-P dumbbell Eqs. (42)–(44) STRidge, a-Lasso Eq. (46) 29
FENE dumbbell N/A STRidge [^8] The results are shown in the supplementary material., a-Lasso Eq .(46) 29
Figure 3: Training data obtained by the UCM model (cf. Eqs. (19)–(21)) (a) under steady shear flow ($\kappa_{xy}={\dot{\gamma}}$) and (b) under oscillatory shear flow ($\kappa_{xy}=\gamma_{0}\omega\cos(\omega t)$). The numbers of total terms obtained (c) from the training data (a) and (d) from the training data (b). (e) The constitutive equations obtained by Rheo-SINDy. In (b), $xx$-, $yy$-, and $xy$-components of the stress tensor are plotted with the black solid, red dotted, and blue dash-dotted lines, respectively. In (c) and (d), the numbers of total terms for five different optimization methods are plotted against the hyperparameter $\alpha$. The black horizontal lines in (c) and (d) indicate the correct number of the terms in the UCM model.
III.3.2 FENE-P Dumbbell Model
@@
$\frac{{\rm d}{\bm{C}}}{{\rm d}t}-{\bm{C}}\cdot{\bm{\kappa}}^{\rm T}-{\bm{\kappa}}\cdot{\bm{C}}=-f_{\rm FENE}(t){\bm{C}}+\frac{n_{\rm K}}{3}{\bm{I}}.$ (35)
The stress tensor is thus obtained by
@@
$\frac{{\rm d}{\bm{\tau}}}{{\rm d}t}-{\bm{\tau}}\cdot{\bm{\kappa}}^{\rm T}-{\bm{\kappa}}\cdot{\bm{\tau}}=-f_{\rm FENE}(t){\bm{\tau}}+2{\bm{D}}+\frac{{\rm D}\ln Z}{{\rm D}t}({\bm{\tau}}+{\bm{I}}),$ (40)
@@
To generate noise-free training data, we solve Eqs. (36)–(39) with $n_{\rm K}=10$ and $\Delta t=1\times 10^{-4}$ for $0\leq t\leq 100$. We apply the oscillatory shear flow with $\gamma_{0}=2$ or $8$ and various $\omega$ values ($\omega\in\{0.1,0.2,\ldots,1\}$). From the computed stress data, we collect data at the interval of $\Delta t_{\rm train}=1\times 10^{-2}$.
III.3.3 FENE Dumbbell Model
@@ IV Results and Discussions
In this section, we present the results of the case studies for five models explained in Sec. III. Through case studies on phenomenological constitutive equations (i.e., the UCM and Giesekus models), we first investigate two key questions for deriving accurate equations: (i) whether oscillatory shear tests or simple (steady) shear tests are more appropriate, and (ii) which of the five optimization methods presented in Sec. II is most suitable. Subsequently, using the experimental and optimization methods identified as appropriate by this investigation, we attempt to obtain the constitutive equations of the dumbbell models. The methods of identifying constitutive models are summarized in Table 3. In the subsequent sections, we examine how the hyperparameter $\alpha$ affects the Rheo-SINDy results. To discuss the effect of the hyperparameter $\alpha$, we also examine the results with various $\alpha$ ($10^{-9}\leq\alpha\leq 10^{3}$). Among the $\alpha$ values examined, we pick $\alpha$ with the smallest number of terms among the results whose error has the same order as the minimum error. For comparison, Rheo-SINDy results with $\alpha$ other than the one chosen by this criterion are also shown.
IV.1 Upper Convected Maxwell Model
Through this case study, we first check the appropriate methods to take the shear rheological data for Rheo-SINDy. Figure 3 shows the training data and results for the UCM model. Figure 3(a) and (b) are the stress data under the steady shear flows with the various shear rates and those under the oscillatory shear flow.
We conducted the Rheo-SINDy regressions by using the polynomial library that includes up to third order terms of $\tau_{xx}$, $\tau_{yy}$, $\tau_{xy}$, and $\kappa_{xy}$. Thus, there were $35$ candidate terms for each component of the constitutive equation. The terms related to $\tau_{zz}$ were excluded because they do not contribute to the UCM dynamics. Figures 3(c) and (d) present the numbers of total terms varying with the hyperparameter $\alpha$ obtained by Rheo-SINDy using the training data (a) and (b), respectively. Figure 3(c) indicates that sparse solutions can be obtained by the STLSQ, STRidge, and a-Lasso, but not by the Lasso and E-Net. Moreover, regarding the number of terms, the STLSQ and STRdge exhibit similar behavior. Specifically, the STLSQ and STRidge with $3\times 10^{-3}\leq\alpha\leq 3\times 10^{-1}$ successfully discovered equations including the correct number of terms. Figure 3(d) indicates that the STLSQ, STRidge, and a-Lasso yielded the correct number of terms, though all five methods gave sparse solutions. In most of the cases where the number of terms obtained was correct, the obtained coefficients were also correct. These results suggest that the oscillatory shear test is more appropriate than the steady shear test to obtain the correct constitutive equations for the UCM model. Figure 3(e) lists the constitutive equations obtained by the STRidge and a-Lasso. We can see that the STRidge and a-Lasso can give the correct constitutive equations, except for the a-Lasso in the steady shear test. Furthermore, we confirmed that the correct equations were obtained even for $\alpha$ values not shown in Fig. 3(e) in the case of the UCM model. These findings show the basic validity of finding the constitutive equations from the rheological data by Rheo-SINDy. Figure 3 indicates that the STLSQ, STRidge, and a-Lasso demonstrate better performance in discovering the correct constitutive equations compared to the Lasso and E-Net; thus, we use the former three methods in the following discussion.
IV.2 Giesekus Model
Figure 4: (a) The number of total terms and (b) the error rate of the constitutive equations obtained by Rheo-SINDy for the Giesekus model. The used optimization methods are the STLSQ (green squares), STRidge (red reverse triangles), and a-Lasso (blue triangles). The regressions were conducted with the multiple ($10$) data trajectories of $\kappa_{xy}=\gamma_{0}\omega\cos(\omega t)$ with $\gamma_{0}=2$ and $\omega\in\{0.1,0.2,\ldots,1\}$ for $0\leq t\leq 100$, respectively.
@@ IV.2 Giesekus Model
Figure 5: (a) The constitutive equations obtained by Rheo-SINDy with (i) the STLSQ, (ii) STRidge, and (iii) a-Lasso from the multiple data trajectories, and (b) their test simulation results under the oscillatory shear flow with $\gamma_{0}=4$ and $\omega=0.5$. The training data are the same as those in Fig. 4. The exact equations for the Giesekus model under shear flow are shown in Eqs. (23)–(26). In (b), the $xx$-, $yy$-, and $xy$-components of the stress tensor are shown with black, blue, and red lines, respectively. The dotted and solid lines in (b) denote the predictions by the equations shown in (a) and those by the exact Giesekus model, respectively.
Figure 6: (a) The total number of terms and (b) the constitutive equations obtained by Rheo-SINDy with the STRidge (black open symbols) and a-Lasso (red closed symbols) for the Hookean dumbbell model (Eqs. (19)–(21)). In (a), the horizontal line indicates the correct number of terms, and circle, triangle, and square symbols represent the results for $N_{\rm p}=10^{3}$, $10^{4}$, and $10^{5}$, respectively.
We here explain the results of Rheo-SINDy for the Giesekus model. This case used the polynomial library consisting of up to second-order terms of $\tau_{xx}$, $\tau_{yy}$, $\tau_{xy}$, and $\kappa_{xy}$, which contain sufficient candidate terms to obtain the exact equations. Figure 4 shows (a) the total number of terms and (b) the error rate obtained by Rheo-SINDy for the training data of the Giesekus model. The error rate is defined as the sum of the mean squared errors (MSEs) of $\dot{\bm{t}}_{\mu\nu}-{\bm{\Theta}}\hat{\bm{\xi}}_{\mu\nu}$. The MSEs were scaled so that the maximum value of each method was 1. Figure 4(a) and (b) indicates that the a-Lasso evidently provides a sparser solution compared to the other two methods when the error rates are comparable. We note that, similar to the number of terms, coefficient values generally depend on $\alpha$.
Figures 5(a) and (b) show the constitutive equations found by Rheo-SINDy and the test simulation results, respectively. For test simulations shown in Fig. 5(b), we employed the oscillatory shear flow with $\gamma_{0}=4$ and $\omega=0.5$, which is outside of the parameters in the training data described in Sec. III.2. Figure 5(a) reveals that the STRidge with $\alpha=3\times 10^{-1}$ gave almost exact constitutive equations, including the value of $\alpha_{\rm G}$ (cf. Eqs. (23)–(26)). As inferred from this, the predictions based on the constitutive equations obtained by the STRidge demonstrate a good agreement with the test data, as shown in Fig. 5(b-ii). In contrast to the success of the STRidge, the STLSQ and a-Lasso failed to identify the correct solution, as indicated in Fig. 5(a). The constitutive equations obtained by the STLSQ with $\alpha=3\times 10^{-1}$ has a low error rate as shown in Fig. 4(b), but its predicted $\tau_{xx}$ significantly deviates from the test data as seen in Fig. 5(b-i). In contrast, although the a-Lasso did not provide the correct solution for $\tau_{xx}$, the test simulations with the obtained constitutive equations exhibit a good agreement with the test data. These test simulations demonstrate that the STRidge and a-Lasso are promising approaches for Rheo-SINDy.
IV.3 Hookean Dumbbell Model
We next explain the results for the Hookean dumbbell model. We here used the polynomial library that includes up to second order terms of $\tau_{xx}$, $\tau_{yy}$, $\tau_{xy}$, and $\kappa_{xy}$. Thus, the total number of candidate terms was $N_{\bm{\Theta}}=15$ for each component.
Figure 6 shows the Rheo-SINDy results for the Hookean dumbbell model with the different numbers of dumbbells. We note that the standard deviation of $\bm{\tau}$ in the training data decreases proportionally with $N_{\rm p}^{-1/2}$. From Fig. 6(a), as the value of $N_{\rm p}$ increases, sparser solutions are obtained, especially for Rheo-SINDy with the STRidge. Unlike the case of the UCM model (cf. Fig. 3), which can be considered as the “noise-free” case of the Hookean dumbbell model, the STRidge provides the correct number of terms only within a narrow range of $\alpha$ values. Nevertheless, if we choose the appropriate $\alpha$ value, the nearly correct constitutive equations can be found by the STRidge, as shown in the upper part of Fig. 6(b). We note that the terms containing $\tau_{yy}$ appear in the time evolution equation for $\tau_{xy}$ obtained by the STRidge. Although these terms do not affect the predictions because ${\tau}_{yy}=0$, we speculate that the appearance of these terms is due to the correlation effects of the noise in $R_{x}$ and $R_{y}$ on the stress (cf. Eq. (31)). When comparing the STRidge and a-Lasso, it is evident that the a-Lasso provides stable and sparse solutions across a broader range of $\alpha$ values, regardless of the $N_{\rm p}$ value. Furthermore, we confirm that the correct equations can be obtained using the a-Lasso, as shown in the lower part of Fig. 6(b). Comparing the results obtained by the a-Lasso with different $N_{\rm p}$, we found that the correct number of terms is obtained in almost the same $\alpha$ range for $N_{\rm p}\geq 10^{4}$. This indicates that using $N_{\rm p}=10^{4}$ provides sufficient results in this case. This partially suggests the effectiveness of the a-Lasso in discovering essential terms from noisy data.
IV.4 FENE-P Dumbbell Model
@@
We next examine whether Rheo-SINDy can find more complex differential equations (i.e., the FENE-P dumbbell model) than the UCM model and the Giesekus model. We utilized Rheo-SINDy with $\bm{T}$ replaced by $\bm{C}$ to discover the differential equations for the conformation tensor $\bm{C}$ of the FENE-P dumbbell model explained in Sec. III.3.2. In this case, we prepared the following custom library:
@@
Figure 7: (a) The total number of terms and (b) the error rate of the constitutive equations obtained by Rheo-SINDy with the STRidge (black squares) and a-Lasso (red reverse triangles) for the FENE-P dumbbell model in the conformation expression (Eqs. (37)–(39) with $n_{\rm K}=10$). The horizontal line in (a) indicates the correct number of terms. The blue circles represent the $\alpha$ values selected for test simulations. Based on our criterion, $\alpha=1\times 10^{-1}$ for the STRidge and $\alpha=3\times 10^{-8}$ for the a-Lasso are picked.
Figure 8: The constitutive equations obtained by Rheo-SINDy with the STRidge and a-Lasso for the FENE-P dumbbell model in the conformation expression (Eqs. (37)–(39) with $n_{\rm K}=10$). Based on our criterion, the appropriate $\alpha$ values were determined as $\alpha=1\times 10^{-1}$ for the STRidge and $\alpha=3\times 10^{-8}$ for the a-Lasso.
Figure 9: Test simulation results under the oscillatory shear flow with $\gamma_{0}=4$ and $\omega=1$ using the equations obtained by Rheo-SINDy with (a) the STRidge and (b) a-Lasso for the FENE-P dumbbell model in the conformation expression. The obtained data from the equations were converted to the stress using the dimensionless form of Eq. (36). The black, blue, and red lines show $\tau_{xx}$, $\tau_{yy}$, and $\tau_{xy}$. The bold, thin dotted, and thin solid lines indicate the exact solutions (Eqs. (36)–(39) with $n_{\rm K}=10$), predictions with smaller $\alpha$ values ($\alpha=1\times 10^{-3}$ for the STRidge and $\alpha=3\times 10^{-8}$ for the a-Lasso), and predictions with larger $\alpha$ values ($\alpha=1\times 10^{-1}$ for the STRidge and $\alpha=1\times 10^{-4}$ for the a-Lasso), respectively.
Figure 10: (a) The total number of terms and (b) the error rate of the constitutive equation obtained by Rheo-SINDy with the STRidge (black squares) and the a-Lasso (red reverse triangles) for the FENE-P dumbbell model in the stress expression (Eqs. (42)–(44) with $n_{\rm K}=10$). The horizontal short-dashed line in (a) indicates that the number of terms is zero. The blue circles represent the $\alpha$ values selected for test simulations. Based on our criterion, the appropriate $\alpha$ values were determined as $\alpha=1\times 10^{-3}$ for the STRidge and $\alpha=1\times 10^{-8}$ for the a-Lasso.
Figure 11: The constitutive equations obtained by Rheo-SINDy with the STRidge and a-Lasso for the FENE-P dumbbell model in the stress expression (Eqs. (42)–(44) with $n_{\rm K}=10$). Based on our criterion, the appropriate $\alpha$ values were determined as $\alpha=1\times 10^{-3}$ for the STRidge and $\alpha=1\times 10^{-8}$ for the a-Lasso.
@@
Figure 7 displays the results: (a) the total number of predicted terms and (b) the error rate as a function of the hyperparameter $\alpha$. The error rate is defined as the sum of the mean squared errors (MSEs) of $\dot{\bm{t}}_{\mu\nu}-{\bm{\Theta}}\hat{\bm{\xi}}_{\mu\nu}$ scaled so that the maximum value of each method is 1. Similar to the results for the phenomenological constitutive equations shown in Secs. IV.1 and IV.2, the a-Lasso provides sparser solutions than the STRidge, and the STRidge gives lower error rates than the a-Lasso. Figure 8 presents the differential equations obtained by the STRidge and a-Lasso for two $\alpha$ values: the one selected by our proposed criterion (i.e., with the smallest number of terms, while the corresponding error is of the same order as the minimum error), and the other has a larger $\alpha$ value (i.e., with a smaller number of terms). From the lower part of Fig. 8, while the a-Lasso provides sparser solutions, they are not correct (cf. Eqs. (37)–(39)). Specifically, in all cases for $C_{xx}$, $C_{yy}$, and $C_{zz}$, the a-Lasso failed to identify the constant term in Eqs. (37) and (38), which is a possible source of larger errors compared to the STRidge. In the case of the STRidge, we confirmed that by choosing $\alpha$ based on our criterion ($\alpha=1\times 10^{-1}$), nearly correct differential equations can be obtained, as shown in the upper part of Fig. 8. Since the $yy$-component and $zz$-component of the stress are equivalent, the exact equations can be recovered by setting $C_{yy}=C_{zz}$. Thus, we conclude that the correct differential equations for the FENE-P dumbbell model can be obtained by Rheo-SINDy.
@@
To validate the obtained equations in Fig. 8, we generated data using the equations under conditions different from those used to generate the training data (namely, the oscillatory shear flow with $\gamma_{0}=4$ and $\omega=1$), and converted these data to the stress data using the dimensionless form of Eq. (36). Figure 9 compares the generated data with the correct data generated from Eqs. (36)–(39). As shown in this figure, the equations obtained by the STRidge can reproduce the exact solutions even when the equations are not exactly correct ($\alpha=1\times 10^{-3}$). In contrast, the test simulations with the differential equations obtained by the a-Lasso show the deviations from the test data, especially for $\tau_{xx}$. In particular, although the equations with $\alpha=3\times 10^{-8}$ have the small error rate for the training data (cf. Fig. 7(b)), those predictions deviate from the test data, and this deviation increases over time. These results emphasize the need to choose an appropriate optimization method to obtain reasonable solutions.
@@
Figure 12: Test simulation results under the oscillatory shear flow with $\gamma_{0}=4$ and $\omega=1$ using the constitutive equations obtained by Rheo-SINDy with (a) the STRidge and (b) a-Lasso for the FENE-P dumbbell model in the stress expression. The black, blue, and red lines represent the $xx$-, $yy$-, and $xy$-components of the stress tensor, respectively. The bold lines show the exact solutions. The thin solid and short-dashed lines indicate the results with smaller $\alpha$ values ($\alpha=1\times 10^{-3}$ for the STRidge and $\alpha=1\times 10^{-8}$ for the a-Lasso) and with larger $\alpha$ values ($\alpha=1$ for the STRidge and $\alpha=1\times 10^{-4}$ for the a-Lasso), respectively.